OpenStudy (girlgoyleh.):

arrange fractions from least to greatest? 3/5 -5/7 2/9

OpenStudy (anonymous):

-5/7 2/9 3/5

OpenStudy (girlgoyleh.):

how'd you get that?

OpenStudy (anonymous):

-5/7 is already negative.. therefore it is the least then for 2/9 and 3/5 you have to do this like that 2/9 = 10/45 and 3/5= 15/45 therefore we have 2/9 less than 3/5 rearranging all, we get -5/7 2/9 3/5 hope u understand

OpenStudy (anonymous):

I absolutely agree with @yajna

OpenStudy (anonymous):

{ 3/5, -5/7, 2/9 }//N -> { 0.6, -0.714286, 0.222222 }